When a number is divided by 10, the remainder is 9. When it is divided by 100, the
remainder is 89. When it is divided by 1000, the remainder is 789. Which of the
following could be that number?
A) 2 987 B) 6 789 C) 2 189 D) 987 E) 1 000

Answer by mananth(16949)

6,789
6789/1000= 6000 +789
..
6789/100 = 6700 +89
..
6789/10 = 6780+9